# A Diagnostic to Find and Help Combat Positivity Issues – with a Focus on Continuous Treatments

### v1

@@ 3.2 Extrapolation and Positivity
When predictions are made using continuous covariates, one typically has to resort to inter- or extrapolation. This can be seen as a sparsity problem as the number of values the continuous covariate can take is higher than what the finite data sample can provide. In this work we consider interpolation as a special case of extrapolation and therefore summarize both under the term extrapolation. As extrapolation is a necessity for the estimation of a causal effect, at least when the treatment or some covariates are continuous and, as discussed before, extrapolation is a possible solution to positivity violations, one always has to consider positivity in conjunction with extrapolation in the continuous treatment case. Especially when considering stochastic positivity violations, the mere concept only makes sense if extrapolation is considered, otherwise finite data on a continuous treatment would always imply a stochastic positivity violation almost everywhere. For theoretical positivity violations, this does not follow automatically. However, we would argue that true theoretical positivity violations are extremely rare from an estimation point of view. If the estimand is defined, i. e., a reasonable scientific question is investigated, it follows that the treatment level of the intervention is – in theory – possible for a given subject. Due to the focus on estimability in this work, we will assume that as given, since the causal road map requires deciding on the target of interest before considering estimability or even identifiability [13]. Once illogical treatments are not a concern anymore, it is rather difficult to end up with a treatment level which is logically possible and at the same time impossible for the data generating process of the observed data set. Even in cases of illegality or contraindication there is a non-zero chance of non-adherence, meaning the true density may be extremely low, but not zero. Therefore, these sources which are typically cited for a deterministic violation [1, 3, 14, 15] are more realistically also just very strong stochastic positivity violations.
@@
As extrapolation is at the core of positivity in continuous treatment cases, the question remains how models extrapolate based on finite data. In a standard linear regression model, every data point potentially influences the estimated line in the same way. For more advanced state-of-the-art models like semi-parametric regression [16] or other cross-validated loss-based learning, e. g. tree-based methods [17], estimation is typically done locally, meaning close data points influence the estimate at a certain point more than data points that are farther away. Therefore, when considering extrapolation, stochastic positivity violations are present in areas where there is not enough data in their surroundings. This leaves two questions open, which a diagnostic that identifies sparsity issues should deal with:
@@ 3.3 The Diagnostic in Simple Terms: A Data-centric Diagnostic
At its core, the diagnostic we propose is a rather intuitive tool. It breaks down the $p$-dimensional information of whether there is data close to the points for which we want to get a prediction into one dimension by using a proximity measure. In the construction of this proximity measure the leading question has to be “How far are we willing to extrapolate information for this specific estimand?” which answers the first question of the last subsection: how do we define close enough? If we are interested in the effect of height on job interview success it makes sense to assume that in order to predict the success of a person of 1.80 m the outcome of a person of 1.81 m is very helpful. If the question concerns whether a person can pass trough a door of 1.80 m without bending down, extrapolation is less successful. While in theory close enough should be decided on using domain knowledge, we realize this is rarely feasible for practitioners. If domain knowledge is not readily available, it might make sense to come up with a best-case, worst-case and median-case scenario. Later on, we also provide some rules of thumb.
Whether a point is close or not should ideally not be a black-and-white issue. Therefore, we propose the use of a “kernel of information” $k_{i}$, which weights points according to the distance of the point we want to make a prediction for, $o_{i}^{int}$, to all values of $o^{obs}$, i. e., to all observed data. The assumption is that the more distant an observation $o_{j}^{obs}$ is from $o_{i}^{int}$ the lower is the information contained within it to help a make a prediction for $o_{i}^{int}$. The kernel $k_{i}$ should fulfill the following properties:
@@
$k_{i}(o^{obs}_{j})\subseteq[0,1]\,\,\forall j\in 1,...,n$
While a completely different kernel $k_{i}$ could be used for every $o_{i}^{int}$, the simplest way of creating $n$ $k_{i}$ is shifting the same kernel considering $o_{i}^{int}$, so it fulfills the first property. The diagnostic measures the effective data points (EDP), where points further away from the point of interest are discounted.
@@
(a) Example of a kernel for a given $o_{i}^{int}$: The values the green kernel has at each data point are summed up and represent the number of effective data points for estimation of outcome $y_{i}^{int}$.
@@
     $k_{i}(x)=k(x+o_{i}^{int})$ (center $k_{i}$ around $o_{i}^{int}$, such that $k_{i}(o_{i}^{int})=1$)
     $\text{EDP}=\sum_{j=1}^{n}\left(k(o_{j}^{obs})\right)$
@@
After the new value $o_{i}^{int}$ is calculated based on $o_{i}^{obs}$ and the intervention scheme $i=0,...,n$, we iterate over the values $o_{i}^{n}$. For each value of $o_{i}^{int}$, the kernel function is evaluated on all $o^{obs}$ and the results are summed up. The $n$ values, one for all values of $o^{int}$ are then returned and can be displayed as a boxplot. This boxplot can then be used in order to answer the second question posed in Section 3.2 in a more nuanced manner. As the lower ends of the boxplot, i. e., the points that are not well supported in the data, are potential problems and therefore more interesting, for this work we modified the typical boxplot to end the whiskers always at the 5th and 95th percentile. In order to illustrate how the diagnostic can be used, we provide a (fictional) example in Figure 1(b). In this plot two estimands are compared based on the diagnostic values for each data point. The plot shows that estimand 1 has more than 25 % of the data completely unsupported, meaning these subjects received a treatment intervention such that they are now in areas of the $p$-dimensional space where no data where observed close by. For estimand 2, most intervened on data points are in areas of high support, however, there are a few data points with support close to zero. In this case, estimand 1 is expected to have a high bias due to sparsity. Estimand 2 will be biased because estimation is bound to perform poorly for a few badly supported data points, however, these data points can be examined and how much impact they might have can be gauged based on domain knowledge. Alternatively, these data points could be excluded from the study via a change in estimand by adjusting the target population.
@@ 3.4 A More Thorough Look: an estimation-focused diagnostic
This simple version of the diagnostic ignores that in the estimation of causal effects, there is not only one model that can potentially be estimated, but two: $\bar{Q}_{n}$ and $g_{n}$, as was discussed in Section 3.1. In order to consider how sparsity influences estimation, we have to examine the two models $\bar{Q}_{n}$ and $g_{n}$, separately. Estimation of $\bar{Q}_{n}$ and $g_{n}$ relies on having observed enough data points $o^{obs}$ close to each value that has to be predicted, i. e. the intervened on observations $o^{int}$. Specifically, for $\bar{Q}_{n}$ that concerns data dimensions $(A,L)$ and for $g_{n}$ only dimensions $L$. If the meaning of close coincides for the two models, i. e., extrapolation can be done similarly for both models, then sparsity for $g_{n}$ implies sparsity for $\bar{Q}_{n}$. As was noted before, if $g_{n}$ is used in estimation, not only bias in $g_{n}$ poses a risk for bias in the overall estimate. If estimated values of $g_{n}$ are small they can amplify bias since the inverse of $g_{n}$ is used as a weight. Therefore, if we want to provide a thorough look at estimability with our diagnostic, the size of the weights $g_{n}$ might produce for estimation needs to be considered as well.
For the estimation-focused diagnostic, the simple data-based diagnostic has to be applied to $(A,L)$ with a kernel describing extrapolation for $\bar{Q}_{n}$, and to $L$, with a kernel describing extrapolation for $g_{n}$. In addition, the diagnostic can be used in order to get a rough idea on the values for the weights $w_{i}=1/g_{n}(o_{i}^{int})$. Since we are interested in a density, the number of points in the surrounding area is of interest as well. For the weights we consider dimensions $(A,L)$ again. As Petersen et al. [3] point out, simply checking the weights for each observation is not enough as there could be regions which do not contain any data points, but if those points were observed they would have extreme weights. Therefore, we will consider the weight of a point exactly where we need it most, i. e., $o_{i}^{int}$. If this has a high weight, it means either there are points here and the high weights make the estimate unstable or there are no points here, but the estimator would place a lot of weight on an observation that was here, meaning the estimate is likely biased by the omission of that point. We suggest using a rather steep kernel as the meaning of the kernel shifts from extrapolation of information from other points to knowledge of the data distribution at this exact point. Instead of the ability to extrapolate, the kernel shape now describes the smoothness of the estimated density.
@@
A diagnostic with these three subdiagnostics could look like the fictional Figure 2. Here estimand 1 indicates severe problems for the estimation of $\bar{Q}_{n}$ and while $g_{n}$ seems to be well estimable from the data, the weights for a lot of data points are extremely high (we chose to set a maximum weight of 1000 for the graphic, so the actual weight is even higher for a lot of points). Therefore, neither g-computation nor IPTW look promising here. For estimand 2, $g_{n}$ is well estimable again, but there are still over 5 % of weights over 1000, so IPTW is not recommended here either. G-computation, however, might be a more promising route to take. While a few points show low data support, most data points are well supported. It might be worthwhile invesigating why some points have so little support and whether a slight change of estimand 2, e. g., by excluding some points form the analysis, might be even more estimable, but even if that is not possible, g-computation while not without flaws, does appear to be a promising approach for estimation.
@@ 3.5 Kernel Choice and Hyperparameters
We recommend using a Gaussian kernel as the default option as it is simple and familiar to most researchers and fulfills all criteria we have stated for kernel. In this section we give some pointers on how the hyperparameters of the kernel can be set, including rules of thumb. First, we recommend thinking about the steepness, in case of the Gaussian the variance, of the kernel in terms of halves. How much would a data point need to differ from $o_{i}^{int}$ in only one dimension, meaning all other features are the same, for the point to contain half as much information as a point which exactly equals $o_{i}^{int}$. Once this question is answered for all dimensions, the kernel hyperparameters can be set accordingly. Making these kinds of assumptions might seem very odd, but we would argue that researchers already make a similar assumption. Whenever a researcher decides to model a curve with $p$ dimensions and $n$ data points, an assumption on the smoothness of the curve and therefore implicitly on how far extrapolation can work is made. While the diagnostic forces the applied researcher to make this assumption very explicit and give hard numbers, the assumption is the same at its core. However, providing exact numbers for the hyperparameters is unlikely to be something a researcher can provide unless they are very well versed in the field of interest. Therefore, we provide some rules of thumb based on $n$, in order to formalize the vague idea that the data are enough to estimate the desired curve in kernel hyperparameters. We recommend one standard deviation as half distance for variables in the adjustment set and half a standard deviation for the treatment effect. Those are the values we used in the practical application in Section 5. Based on this rule-of-thumb, one could formulate a worst case scenario in order to get more insight on how dependent on these hyperparameters the results are. We recommend to halve these values for the worst case scenario leading to lower faith in extrapolation in the data.
In very high-dimensional cases it might be impossible to find data points which are close to $o_{i}^{int}$ in every dimension. In these cases it might make sense to use the harmonic mean of the kernel value in every single dimension or set a minimum value for each dimension and then multiply the single dimension values. Otherwise the diagnostic might be overly pessimistic in cases where a large difference in a single dimension does not influence estimability a lot.

### v3

@@ 3.2 Extrapolation and Stochastic Positivity
When predictions are made using continuous covariates, one typically has to resort to inter- or extrapolation. This can be seen as a sparsity problem as the number of values the continuous covariate can take is higher than what the finite data sample can provide. In this work we consider interpolation as a special case of extrapolation and therefore summarize both under the term extrapolation. As extrapolation is a necessity for the estimation of a causal effect, at least when the treatment or some covariates are continuous and, as discussed before, extrapolation is a possible solution to stochastic positivity violations, one always has to consider stochastic positivity in conjunction with extrapolation in the continuous treatment case. In fact, the mere concept of a stochastic positivity violation only makes sense if extrapolation is considered, otherwise finite data on a continuous treatment would always imply a stochastic positivity violation almost everywhere.
@@
As extrapolation is at the core of stochastic positivity in continuous treatment cases, the question remains how models extrapolate based on finite data. In a standard linear regression model, every data point potentially influences the estimated line in the same way. For more advanced state-of-the-art models like semi-parametric regression [17] or other cross-validated loss-based learning, e. g. tree-based methods [18], estimation is typically done locally, meaning close data points influence the estimate at a certain point more than data points that are farther away. Therefore, when considering extrapolation, stochastic positivity violations are present in areas where there is not enough data in their surroundings. This leaves two questions open, which a diagnostic that identifies sparsity issues should deal with:
@@ 3.3 The Diagnostic in Simple Terms: A Data-centric Diagnostic
At its core, the diagnostic we propose is a rather intuitive tool. It breaks down the $p$-dimensional information of whether there is data close to the points for which we want to get a prediction into one dimension by using a proximity measure. In the construction of this proximity measure the leading question has to be “How far are we willing to extrapolate information for this specific estimand?” which corresponds to the first question of the last subsection: “How do we define close enough?”. If we are interested in the effect of height on job interview success it makes sense to assume that in order to predict the success of a person of 1.80 m the outcome of a person of 1.81 m is very helpful. If the question concerns whether a person can pass trough a door of 1.80 m without bending down, extrapolation is less successful. While in theory close enough should be decided on using domain knowledge, we realize this is rarely feasible for practitioners. If domain knowledge is not readily available, it might make sense to come up with a best-case, worst-case and median-case scenario. Later on, we also provide some rules of thumb.
Whether a point is close or not should ideally not be a black-and-white issue. Therefore, we propose the use of a “kernel of information” $k$, which weights points according to the distance of the point we want to make a prediction for, $o_{i}^{int}$, to all values of $o^{obs}$, i. e., to all observed data. The assumption is that the more distant an observation $o_{j}^{obs}$ is from $o_{i}^{int}$ the lower is the information contained within it to help make a prediction for $o_{i}^{int}$. The kernel $k$ is most easily formulated as taking the distance between $o_{i}^{int}$ and $o_{j}^{obs}$, as the kernel needs to be centered around $o_{i}^{int}$, and therefore should fulfill the following properties:
@@
$k(o_{j}^{obs}-o_{i}^{int})\subseteq[0,1]\,\,\forall i,j\in 1,...,n$
The diagnostic measures the effective data points (EDP), where points further away from the point of interest are discounted.
@@
(a) Example of a kernel $k(x)=k(x-o_{i}^{int})$ for a given $o_{i}^{int}$: The values the green kernel has at each data point are summed up and represent the number of effective data points for estimation of outcome $y_{i}^{int}$.
@@
  $\text{EDP}=\sum_{j=1}^{n}\left(k(o_{j}^{obs}-o_{i}^{int})\right)$
@@
After the new value $o_{i}^{int}$ is calculated based on $o_{j}^{obs}$ and the intervention scheme for j=0,,nj=0,...,n, we iterate over the values $o_{i}^{int}$ for i=0,,ni=0,...,n. For each value of $o_{i}^{int}$, the kernel function is evaluated on all $o_{j}^{obs}$ and the results are summed up. The $n$ values, one for all values of $o^{int}$, are then returned and can be displayed as a boxplot. This boxplot can then be used in order to answer the second question posed in Section 3.2 in a more nuanced manner. The lower ends of the boxplot, i. e., the points that are not well supported in the data, potentially indicate a problem and are therefore more interesting. Therefore, we modified the typical boxplot to end the whiskers always at the 5th and 95th percentile for this work. In order to illustrate how the diagnostic can be used, we provide a (fictional) example in Figure 1(b). In this plot, two estimands are compared based on the diagnostic values for each data point. The plot shows that estimand 1 has more than 25 % of the data completely unsupported, meaning these subjects received a treatment intervention such that they are now in areas of the $p$-dimensional space where no data were observed close by. For estimand 2, most intervened on data points are in areas of high support, however, there are a few data points with support close to zero. In this case, estimand 1 is expected to have a high bias due to sparsity. Estimand 2 will be biased because estimation is bound to perform poorly for a few badly supported data points, however, these data points can be examined and how much impact they might have can be gauged based on domain knowledge. Alternatively, these data points could be excluded from the study via a change in estimand by adjusting the target population.
@@ 3.4 A More Thorough Look: An Estimation-focused Diagnostic
This simple version of the diagnostic ignores that in the estimation of causal effects, there is not one model that can potentially be estimated, but two: $\bar{Q}_{n}$ and $g_{n}$, as discussed in Section 3.1. In order to consider how sparsity influences estimation, we have to examine the two models $\bar{Q}_{n}$ and $g_{n}$, separately. Estimation of $\bar{Q}_{n}$ and $g_{n}$ relies on having observed enough data points $o^{obs}$ close to each value that has to be predicted, i. e. the intervened on observations $o^{int}$. Specifically, for $\bar{Q}_{n}$ that concerns data dimensions $(A,L)$ and for $g_{n}$ only dimensions $L$. If the meaning of close coincides for the two models, i. e., extrapolation can be done similarly for both models, then sparsity for $g_{n}$ implies sparsity for $\bar{Q}_{n}$. As was noted before, if $g_{n}$ is used in estimation, it is not only bias in $g_{n}$ that poses a risk for bias in the overall estimate. If estimated values of $g_{n}$ are small they can amplify bias since the inverse of $g_{n}$ is used as a weight. Therefore, if we want to provide a thorough look at estimability with our diagnostic, the size of the weights $g_{n}$ might produce for estimation needs to be considered as well.
For the estimation-focused diagnostic, the simple data-based diagnostic has to be applied to $(A,L)$ with a kernel describing extrapolation for $\bar{Q}_{n}$, and to $L$, with a kernel describing extrapolation for $g_{n}$. In addition, the diagnostic can be used in order to get a rough idea on the values for the weights $w_{i}=1/g_{n}(o_{i}^{obs})$, as the weights are based on $g_{n}$. Checking the weights for extreme values is recommended in literature in order to screen for positivity issues, specifically bias number 4, but as Petersen et al. [4] point out, only considering the weights which are to be used in the model is not enough as there could be regions which do not contain any data points. Some of these data points could, however, have a high weight had they been included in the estimation and been an indication of a sparsity issue. Therefore, we recommend checking the weights of ideal data points, which lie exactly at $o_{i}^{int}$, resulting in $w_{i}^{ideal}=1/g_{n}(o_{i}^{int})$. Those weights can be roughly gauged using a kernel with the idea of counting effective data points at $o_{i}^{int}$. If there are very few observed data points, the inverse of the amount of observed data points is going to become very large, which is an indicator for sparsity inflating the bias. We suggest using a rather steep kernel as the meaning of the kernel shifts from extrapolation of information from other points to knowledge of the data distribution at this exact point. Instead of the ability to extrapolate, the kernel shape now describes the smoothness of the estimated density.
@@
A diagnostic with these three subdiagnostics could look like the fictional Figure 2. Here, estimand 1 indicates severe problems for the estimation of $\bar{Q}_{n}$ and while $g_{n}$ seems to be well estimable from the data, the weights for a lot of data points are extremely high (we chose to set a maximum weight of 1000 for the graphic, so the actual weight is even higher for a lot of points). Therefore, neither g-computation nor IPTW look promising here. For estimand 2, $g_{n}$ is well estimable again, but there are still over 5 % of weights over 1000, so IPTW is not recommended here either. G-computation, however, might be a more promising route to take. While a few points show low data support, most data points are well supported. It might be worthwhile investigating why some points have so little support and whether a slight change of estimand 2, e. g., by excluding some observations from the analysis known as restricting the population (more on that in Section 5.2.1, where we perform a similar adjustment), might be even more estimable, but even if that is not possible, g-computation does appear to be a promising approach for estimation based on the EDP.
@@ 3.5 Kernel Choice and Hyperparameters
In this section we provide some guidance on the choice of kernel, discuss how categorical variables can be dealt with, and present some rules of thumb on how to decide on hyperparameters.
The kernel represents how information is localized in the data and should be chosen in accordance with how extrapolation is expected to work in the data. Therefore, it is related to the model choice, which has to be based on similar considerations. For example, if the true function that is to be estimated is rather smooth, the kernel can be wide and a potential Generalized Additive Model estimator can use fewer basis functions. Many estimators use kernel approaches to assign more weight to data points that are close to the point of interest, leveraging the simple idea that nearby observations are typically more informative than distant ones. Kernel regression for example may be based on $\widehat{\mathrm{E}}(y|x)=\sum_{i=1}^{n}\frac{K_{h}(x_{i}-x)y_{i}}{\sum_{i=1}^{n}K_{h}(x_{i}-x)}$, with kernel function $K(t)=\frac{1}{h}K(\frac{t}{h})$ where $h$ is the bandwidth, in order to model a relationship of independent variable $x$ to dependent variable $y$ in a data set with sample size $n$ [19]. Here, the quantity $nh$ is often referred to as the effective sample size as it describes the amount of data which is locally used for estimation [20]. This quantity is rather similar to the idea of effective data points presented in this paper, as it is a local measure of information in the data. While the effective sample size refers to the information used in estimation, our effective data points aim at gauging the amount of information that is present in the data, which could theoretically be used by an estimator. The bandwidth of the kernel $h$ is only a descriptive measure of how extrapolation is conducted in the estimator, whereas the diagnostic hyperparameters, which control the steepness of the kernel, aim to describe how extrapolation should ideally be performed.
In the simplest case, the kernel function could be a uniform kernel
$k(o_{j}^{\text{obs}}-o_{i}^{int})=\begin{cases}1,&\text{if }\|o_{pj}^{obs}-o_{pi}^{int}\|\leq h_{p}\,\forall p\\ 0,&\text{otherwise,}\end{cases}$
where p=1,,Pp=1,...,P are the dimensions of the data. This reflects a binary choice sorting data points into the categories of useful for estimation, which corresponds to assigning a value of 1, or not useful for estimation, assigning a value of 0. The hyperparameters $h_{p}$ in this case decide where the cutoff between those categories lies. This results in a highly interpretable diagnostic, which shows for each observation how many data points are within that predefined range. Regardless of the kernel in use, there are two special cases of hyperparameter choices which can be expressed in accordance with the formula stated above as $h_{p}=0$ and $h_{p}\to\infty$ for all $p$. If $h_{p}=0$, the assumption is that one can only learn from data points that are exactly alike. Therefore EDP will show how many data points there are with the exact same values in all covariates. If $h_{p}\to\infty$, the assumption is that all data points carry the same amount of information independent of any (dis-)similarities. This assumption is useful for example in a linear case. If the true shape of the function to fit is linear – and as a consequence a linear estimator is used in order to capture the truth as well as possible – the similarity aspect becomes unimportant and all data points should be counted with a weight of one. However, this will lead to $EDP=n$, which is not a useful result. In this case, the diagnostic becomes obsolete.
In causal inference, the treatment and confounders taken together are typically not one-dimensional. Therefore a multivariate kernel is needed. A simple choice would be the Multivariate Gaussian, but other kernel types could be used as well. In particular, univariate kernels could be used to construct multivariate ones, which can be done, for example, using a product kernel or a radial basis function. For a product kernel the $P$ univariate kernels $k_{p}$, one for each dimension of $x$, can be used to create $k(x)=\prod_{p=1}^{P}k_{p}(x_{p})$. This is especially helpful if different types of kernels should be used for different variables, but also comes with the drawback that a specific kernel has to be designed fo every variable. The alternative, the radial basis function $k(x)=k(\|x\|)$, with distance measure $\|.\|$, may be simpler to define. For Gaussian kernels, the two methods are equivalent.
The diagnostic was developed especially with continuous variables in mind. However, it can be used with categorical variables as well. Often, categorical variables will be dummy coded and therefore binary. As a consequence, the kernel (only considering this single dimension) should return $1$ if the data points fall in the same category and can be set to some fixed value $c$ if they fall into different categories. Typically, we would recommend $c=0$ as it is unclear how this categorical feature will influence the outcome in connection with the other variables. However, in cases where the influence of the category is expected to be of lesser importance, a value >0>0 could be chosen, but the resulting EDP will have to be more carefully considered. For example, in a data set with the binary category of sex and using $c=0.5$, an EDP of 20, which might be considered satisfactory depending on the specific context, this could mean that there are 40 similar men but not one woman even though the intention is to predict for a woman in this hypothetical.
@@
The considerations above highlight how difficult the choice of the kernel function is. Ideally, the kernel and its hyperparameters are chosen using expert knowledge on the true shape of the function to be estimated. However, we recognize that this knowledge is hard to come by and we provide some guidance in the following. With regard to the kernel choice, we recommend a Gaussian kernel is used as it is simple, familiar to most researchers, and fulfills the criteria for the kernel function stated above. This leaves the choice of the hyperparameters, which control how steep the kernel is and therefore how far one expects to be able to extrapolate. First, we recommend thinking about the steepness – the variance for the Gaussian case – of the kernel in terms of halves. How much would a data point need to differ from $o_{i}^{int}$ in only one dimension, meaning all other features are the same, for the point to contain half as much information as a point which exactly equals $o_{i}^{int}$. Once this question is answered for all dimensions, the kernel hyperparameters can be set accordingly. Making these kinds of assumptions might seem very odd, but we would argue that researchers already make a similar assumption. Whenever a researcher decides to model a curve with $P$ dimensions and $n$ data points, an assumption on the smoothness of the curve and therefore implicitly on how far extrapolation can work is made. While the diagnostic forces the applied researcher to make this assumption very explicit and provide hard numbers, the assumption is the same at its core. Nevertheless, we recognize that deciding on exact numbers for the hyperparameters is quite challenging even for a researcher well versed in the field in question. Therefore, we provide some rules of thumb based on the observed data. We recommend one standard deviation as half distance for variables in the adjustment set and half a standard deviation for the treatment effect, as this is the most important feature in the prediction model. Those are the values we used in the practical application in Section 5 and while the simulation in Section 4 does not use exactly these values (as one fixed hyperparameter was chosen for several data distributions in order to simplify interpretation), the values used here are similar. In both cases the rule of thumb provided satisfactory results. In case more expert knowledge is available, one could start from these basic rules of thumb and tweak the values based on this knowledge. In addition, one could formulate worst and best case scenarios in order to gain more insight into how dependent on these hyperparameters the results are. We recommend to halve the rule of thumb values provided for the worst case scenario and double them for the best case scenario. Rather than basing the hyperparameters on the standard deviation, which is not robust to outliers, one could base this choice on the median absolute deviation or the interquartile range. Alternatively, one could use the average pairwise distance of two points, which is harder to compute, but provides more information on distances in the data set which are at the core of the diagnostic and therefore might be more interpretable.
In very high-dimensional cases it might be impossible to find data points which are close to $o_{i}^{int}$ in every dimension. This is how the curse of dimensionality is reflected in the EDP. We provide a simulation study illustrating how the diagnostic behaves in higher dimensions as well as some considerations on how to combat these issues in Appendix B.